how many green cars... for every 100 cars made at a factory, 1/2 are red, 20 are blue, and the rest are green.
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


1/2 of 100 = 50 cars are red
20 are blue
50+20 = 70 are red or blue (assume a car can't be both colors at once)
100-70 = 30 cars are green

Answer by ikleyn(52781)

.

We can not say, how many cars are green.

All that we can deduce from the problem, is that for every 100 cars, 30 cars are green.


Also,  when you formulate the problem for the forum,  try to place the info in logical order,
and write your question  IN  FULL  at the end.